OpenStudy (anonymous):

state the geometrical relationship between graphs of any linear function f(x) and its coressponding function g(x) found by reversing the coordinates in the ordered pairs. explain this relationship for varying gradients, discuss whether f(x) and g(x) will not intersect

OpenStudy (thomas5267):

g(x) will be f(x) reflected across the y=x axis.

OpenStudy (anonymous):

anyone?

OpenStudy (loser66):

I don't get it!! it says "any linear function f(x)" and "g(x) = reversing the coordinates in the ordered pairs" If I set f(x) = x, then y =x, hence g(x) is f(x) itself. This is an "any linear" but then f(x) = g(x) , then intersect at many infinite points.

OpenStudy (er.mohd.amir):

if function is linear then its f(x)=ax+b form and reversing it g(x)=(x-b)/a

OpenStudy (loser66):

Unfortunately, it is "Any" , so that if I can give out a counter example, as long as it is a linear function, I am ok.

OpenStudy (loser66):

Yes, linear function has general form y = ax +b. Who say y =x is not a linear function?

OpenStudy (er.mohd.amir):

f(x,y)=ax+by g(y,x)=ay+bx

OpenStudy (er.mohd.amir):

if f is ordered pair .

OpenStudy (anonymous):

f(x)=2x-5 g(x)=x/2+5/2

OpenStudy (anonymous):

does that help?

OpenStudy (thomas5267):

No. The statement is equivalent to that \(g(x)=f^{-1}(x)\). So if \(f(x)=y\), then \(g(y)=x\).

OpenStudy (anonymous):

kk so by inspecting these graphs state the geometrical relationship . explain this relationship. for varying gradients, discuss whether f(x) or g(x) will or will not intersect

OpenStudy (anonymous):

hope this helps

OpenStudy (dan815):



OpenStudy (dan815):